# Mapping and Scheduling Spiking Neural Networks On Segmented Ladder Bus Architectures

### v1

@@ Abstract
Large-scale neuromorphic architectures consist of computing tiles that communicate spikes using a shared interconnect. The communication patterns in such systems are inherently sparse, asynchronous, and localized due to the spiking nature of neural events, characterized by temporal sparsity with occasional high bursts of traffic. These characteristics necessitate interconnects optimized for handling high-activity bursts while consuming minimal power during idle periods. Dynamic segmented bus has been proposed as one of the promising interconnects for such system for its simplicity, scalability and low power consumption. However, deploying spiking neural network applications on such buses presents challenges, including substantial inter-cluster traffic, which can lead to elevated network congestion and spike loss, as well as unnecessary energy expenditure. In this paper, we propose a three step process to deploy spiking neural network applications on dynamic segmented bus that aims to reduce spike loss and conserve energy. Firstly, we formulate optimization heuristics to mitigate spike loss and energy consumption based on application network connectivity. Secondly, we analyze the application traffic to determine spike schedule that minimize traffic flooding. Lastly, we propose a routing algorithm to minimize spike traffic path crossings. We then evaluate our approach using a cycle-accurate network simulator. The simulation results show that our mapping and scheduling algorithms can completely eliminate spike loss while keeping energy consumption much lower compared to conventional NoCs.
@@
journal: arXiv
@@ 1 Introduction
An event-driven neuromorphic system is a computing platform that implements biological neurons and synapses in hardware to execute spiking neural networks (SNN) [1]. To address design scalability, a neuromorphic system is designed as a many-core hardware, where cores are interconnected using a shared time-multiplexed interconnect such as shared bus, network-on-chip (NoC) (e.g., NorthPole [2], DYNAP-SE2 [3]), and segmented bus (e.g., $\mu$Brain [4]).
@@
The traffic in such systems is driven by spiking events, differing fundamentally from the continuous and synchronous communication typical of traditional computing architectures. A defining feature of neuromorphic traffic is temporal sparsity [5, 6]. Communication occurs only when a neuron generates a spike, resulting in sparse data traffic. In large-scale neuromorphic systems, densely connected neurons are typically grouped within the same clusters, while connections between clusters remain sparse. This clustering leads to traffic patterns where inter-cluster communication can experience short periods of intense activity, especially when many neurons spike simultaneously, followed by long periods of inactivity. Consequently, the communication infrastructure supporting neuromorphic system must be optimized to accommodate this characteristic: it must not only handle high-burst traffic efficiently but also consume minimal power during idle states.
@@
Similar to other many-core architectures, SNN applications are required to be mapped to the hardware prior to execution. This process usually includes, but not limited to: partitioning the application into several clusters which can fit into the hardware neural core; placing the clusters into hardware based on the application connectivity; and scheduling/routing the spike traffic between cores. Previous works have been proposed to address mapping SNNs into neuromorphic hardware, such as SpiNeMap [7] and [8]. These works show that optimized mapping can greatly influence the performance of the running SNN application, including energy consumption, latency, and inter-spike interval.
@@
However, these methods mainly perform such mapping operations on NoCs, which is not fully suitable for running SNN applications. The rationale for this can be attributed to the significant energy and area overhead of NoC, which is mainly due to the use of large buffers to temporarily store data packets and look-up tables (LUTs) to store routing tables inside each switch. While effective for conventional data handling, such design becomes resource-intensive in neuromorphic contexts where communication is sparse and irregular. A major factor contributing to this inefficiency is the leakage power associated with the substantial memory capacity required by the buffers. This results in unnecessary energy consumption, which contradicts the low-power and burst-like spike traffic requirements that are critical for neuromorphic systems. To mitigate such disadvantages, bufferless NoCs have been proposed as an alternative to buffered NoCs [9]. However, in bufferless NoCs, when multiple packets compete for the same output port, all but one packet must be deflected and rerouted through alternative paths. This deflection process leads to additional traversals through the network [10], thereby increasing the overall network activity and energy consumption.
@@
Segmented bus has been introduced as an alternative for NoC in these multi-core neuromorphic hardware systems. While the switches in a segmented bus are also bufferless, the communication system is structured to minimize path crossings, effectively avoiding the deflection and rerouting issues typically encountered in bufferless NoCs. It was shown that segmented buses such as NeuSB [11] outperforms NoCs in both latency and energy consumption for running SNN applications. Nevertheless, NeuSB lacks runtime switch controls and thus does not truly support running different applications on the same hardware. In this work, we focus on segmented ladder bus [12], the latest dynamic segmented bus interconnect that can offer runtime configuration for routing and scheduling network traffic.
@@ 2 Background and Related Works
This section presents the requisite background to comprehend the segmented ladder bus [12], the primary interconnect architecture targeted by our mapping and scheduling methodologies. Furthermore, we review related works on the mapping of SNNs to neuromorphic hardware architectures, emphasizing the associated challenges and recent innovations in this domain.
@@ 2.1 Segmented Ladder Bus
A segmented ladder bus is governed globally by a compile time software framework and locally use coarse-grain runtime hardware controllers in order to enable dynamic configuration of switches. These local controllers encompass predefined switch change scenarios and facilitate the transmission of control signals to the switches accordingly. By issuing the control signals from the controllers with the correct timing, the switches can be used to create paths that support multiple simultaneous connections inside the network, matching application requirements. The controllers are implemented in a distributed manner, where each controller is responsible for managing the switches within its vicinity, similar to a distributed loop controller [13]. As a result, in a large-scale network, numerous controllers will be deployed to effectively govern the switches throughout the system.
@@
A more comprehensive understanding of the advantages offered by segmented ladder bus can be gained through an examination of Figure 2. Using the cluster communication graph with a suitable mapping, multiple communication paths can be facilitated simultaneously: Tile 0 sends spikes to Tiles 5 and 6, Tile 1 sends to Tile 2, and Tile 7 to Tiles 3 and 8.
@@ 2.2 Related Works
One prominent method is SpiNeMap [7], which focuses on optimizing the mapping of spiking neural networks (SNNs) onto neuromorphic hardware. SpiNeMap employs a multi-objective optimization strategy to balance energy consumption, latency, and inter-spike interval. By considering the unique communication patterns and computational requirements of SNNs, SpiNeMap achieves significant improvements in the overall performance of neuromorphic systems. Similarly, SNEAP [14] suggested using one of the three heuristics, namely Simulated Annealing, Particle Swarm Optimization, and Tabu search for mapping neuromorphic clusters to NoC. However, their reliance on static NoC topologies limits its scalability and adaptability to dynamic workloads.
Another approach by Jin et al. [8] presents a heuristic-based mapping algorithm that aims to minimize communication overhead in NoC-based neuromorphic systems. It solves the initial placement using the Hilbert curve, a space-filling curve with unique properties that are particularly suitable for mapping SNNs. Afterwards, a Force Directed algorithm is developed to optimize the initial placement. This method leverages the inherent sparsity of SNNs to reduce the number of active communication channels, thereby lowering energy consumption and improving latency. Despite its effectiveness, the heuristic nature of this algorithm may not always guarantee optimal mappings, especially in larger and more complex networks.
NeuMap [15] introduces a machine learning-based technique for mapping neuromorphic clusters on NoC architectures. By training a neural network to predict optimal mappings based on historical data, NeuMap can dynamically adjust to varying workloads and network conditions. This adaptive approach demonstrates significant potential in enhancing the efficiency of neuromorphic systems. Nonetheless, the training overhead and the need for extensive historical data may pose challenges in real-time applications.
@@ 3 Proposed Solutions
In our previous work [16], we have described the SNN application partitioning process in more details. This allocates neurons to memristive crossbar clusters while maximizing the crossbar utilization and minimizing inter-cluster communication. Although such process is essential, it works independently of the types of interconnect hardware. For this reason, it is not the focus of this paper.
Once an SNN application has been partitioned into clusters, we need to: map the clusters to actual hardware tiles; schedule spike traffic; and route the scheduled traffic. Each of these steps can be optimized to improve the hardware utilization of the network, guarantee performance of the application, and reduce energy consumption. Figure 3 shows our proposed solution processes integrated inside our evaluation framework. While numerous state-of-the-art network simulators provide support for Network on Chip architectures, none currently offer built-in capabilities for simulating segmented bus architectures. Thus, we developed a custom network simulator tailored to segmented busesbased on Noxim [17]. Noxim was chosen as the foundation for our simulator due to its open-source, extensible codebase, SystemC-based [18] cycle-accurate simulation engine, and lightweight, scalable design.
@@ 3.1 Cluster Mapping
[…]
Different mappings will have different impacts on energy consumption, latency, and the number of crossed communication paths. In our segmented ladder bus design, communication energy consumption is proportional to the length of the communication paths and the number of switch configurations that need to be changed. Furthermore, unlike NoCs, the bufferless nature of the switches in segmented bus causes simultaneous intersecting communication paths to result in either packet drops or additional delays at the source nodes to avoid packet loss. Thus, we propose two different mapping cost functions specifically adapted to ladder segmented bus to minimize either dynamic energy consumption or number of crossed paths. They are computed as follows.
@@
Given that cluster mapping is a well-established NP-hard problem [19], a practical solution is implemented using a Hill Climbing approach [20]. This metaheuristic provides an efficient and practical way to explored the vast solution space while balancing computational cost and solution quality. We employ a specific variant, called steepest ascent Hill Climbing, to enhance the search process. The algorithm starts with a random mapping solution and compares the cost of that solution with all neighboring solutions. A neighbor is defined as a mapping obtained by swapping the positions of two nodes in the current mapping. This ensures that the search space of the algorithm is limited to only valid mappings, which improve the efficiency of the search. The algorithm moves the solution to the neighbor with the most improved cost. The search continues until there is no neighbor with improved cost. The local search ends here, with the algorithm keeping record of this solution. It is important to note that Hill Climbing may converge to a local minima. To mitigate this, when the search terminates due to the lack of better-cost neighbors, we incorporate a number of random swaps and re-initiate the search procedure. Ultimately, the algorithm concludes its execution after running a predetermined number of random perturbations.
[…]
@@ 6       $k=0$;
7       while $improved$ and $k<I$ do /* while cost is improved and under the number of iteration limit */
@@ 3.2 Spike Traffic Scheduling
With the cluster mapping acquired from algorithm 1 and the spike traffic from CARLSim [21] or SNNTorch [22], we can obtain the exact timing of the spike traffic among the hardware tiles. For any given time with multiple simultaneous communication links, the scheduler would need to find groups of links for which no two links in any group intersect, while minimizing the number of groups to prevent unnecessary delays.
This problem is essentially a bin packing problem with conflicts (BPPC) [23], which is NP-complete. Therefore, instead of seeking an optimal schedule, we develop a custom heuristic to get a solution within a practical computational limit. An example of a spike traffic schedule can be found in Figure 5.
@@ 8                   $crossed=False$;
10                         if $check\_cross(G_{ladder},L,L_{G})$ then /* check whether the current link L crosses with $L_{G}$/
@@ 31             $max\_spike=0$;
    For each link $l$, the algorithm checks existing groups to see if the link can be added without conflict. For each group, the algorithm checks if the new link crosses with any existing link in the group. This involves iterating over each link in the group, and for each pair of links, the check_cross function is called, which can be simplified to take constant time $O(1)$ due to the defined and regular topology of the network. In the worst case, each link will be compared to all other $L-1$ links, giving $O(L^{2})$ comparisons.
@@ 3.3 Path Routing
Our solution to the path routing problem primarily adopts a greedy algorithm, supplemented by A* [24] sub-processes for efficient shortest path finding. We sort all the simultaneous communication links $CLinks$ according to the following order: we prioritize all the links with source and destination tiles on the same bus lane first, followed by sorting based on the orientation of the links, and finally by the number of spikes on each link. The reasoning of this sort is three folds: Firstly, we reduce the number of potential crossing between links by routing the links with source and destination tiles on the same bus lane and the shorter distance ones first. This is because the shortest path for tiles on the same bus lane is unique, while there can be many paths with the same length for tiles on different bus lanes. Secondly, links with the same orientation can be easily routed one by one so that they do not cross each other. Thirdly, by prioritizing to route the links with more spikes, the paths for these links will be shorter, which means lowering the number of spikes that need to travel long distances, which in turn helps in reducing the energy consumption in the network. We then represent the segmented ladder bus interconnect as an undirected graph $G_{ladder}$ with tiles and switches as vertices and bus wires as edges. The initial weights of all vertices and edges in the graph are set to 1. The algorithm processes the sorted communication links sequentially, and uses a modified A* algorithm to determine the shortest weighted path for each specific link. After determining the shortest path for the selected link, the weights of all vertices are increased proportionally to the number of spikes communicated on this particular link. Subsequently, the routed link is removed from the list of communication links. The algorithm continues the process iteratively, selecting and routing the most communicated links one by one, until all paths are effectively routed.
Input: $G_{ladder}$, $CLinks$
Output: $RoutedList$

### v2

@@ Abstract
Large-scale neuromorphic architectures consist of computing tiles that communicate spikes using a shared interconnect. The communication patterns in such systems are inherently sparse, asynchronous, and localized due to the spiking nature of neural events, characterized by temporal sparsity with occasional bursts of traffic. These characteristics necessitate interconnects optimized for handling high-activity bursts while consuming minimal power during idle periods. Dynamic segmented bus has been proposed a promising interconnect for its simplicity, scalability and low power consumption. However, deploying spiking neural network applications on such buses presents challenges, including substantial inter-cluster traffic, which can lead to network congestion, spike loss, and unnecessary energy expenditure. In this paper, we propose a three-step process to deploy SNN applications on dynamic segmented buses aiming to reduce spike loss and conserve energy. Firstly, we formulate optimization heuristics to mitigate spike loss and energy consumption based on application connectivity. Secondly, we analyze the application traffic to determine spike schedules that minimize traffic flooding. Lastly, we propose a routing algorithm to minimize spike traffic path crossings. We evaluate our approach using a cycle-accurate network simulator. The simulation results show that our algorithms can eliminate spike loss while keeping energy consumption significantly lower compared to conventional NoCs.
@@
journal: Journal of Systems Architecture
@@ 1 Introduction
[…]
@@
The traffic in such systems is driven by spiking events, differing fundamentally from the continuous and synchronous communication typical of traditional computing architectures. A defining feature of neuromorphic traffic is temporal sparsity [olshausen2004sparse, wolfe2010sparse]. Communication occurs only when a neuron generates a spike, resulting in sparse data traffic. In large-scale neuromorphic systems, densely connected neurons are typically grouped within the same clusters, while connections between clusters remain sparse. This clustering leads to traffic patterns where inter-cluster communication can experience short periods of intense activity, especially when many neurons spike simultaneously, followed by long periods of inactivity. Consequently, the communication infrastructure supporting neuromorphic system must be optimized to accommodate this characteristic: it must not only handle high-burst traffic efficiently but also consume minimal power during idle states.
@@
Similar to other many-core architectures, SNN applications are required to be mapped to the hardware prior to execution. This process usually includes, but not limited to: partitioning the application into several clusters which can fit into the hardware neural core; placing the clusters into hardware based on the application connectivity; and scheduling/routing the spike traffic between cores. Previous works have been proposed to address mapping SNNs into neuromorphic hardware, such as SpiNeMap [spinemap] and [jin2023mapping]. These works show that optimized mapping can greatly influence the performance of the running SNN application, including energy consumption, latency, and inter-spike interval.
@@
However, these methods mainly perform such mapping operations on NoCs, which is not fully suitable for running SNN applications. The rationale for this can be attributed to the significant energy and area overhead of NoC, which is mainly due to the use of large buffers to temporarily store data packets and look-up tables (LUTs) to store routing tables inside each switch. While effective for conventional data handling, such design becomes resource-intensive in neuromorphic contexts where communication is sparse and irregular. A major factor contributing to this inefficiency is the leakage power associated with the substantial memory capacity required by the buffers. This results in unnecessary energy consumption, which contradicts the low-power and burst-like spike traffic requirements that are critical for neuromorphic systems. To mitigate such disadvantages, bufferless NoCs have been proposed as an alternative to buffered NoCs [moscibroda2009case]. However, in bufferless NoCs, when multiple packets compete for the same output port, all but one packet must be deflected and rerouted through alternative paths. This deflection process leads to additional traversals through the network [michelogiannakis2010evaluating], thereby increasing the overall network activity and energy consumption.
@@
Segmented bus has been introduced as an alternative for NoC in these multi-core neuromorphic hardware systems. While the switches in a segmented bus are also bufferless, the communication system is structured to minimize path crossings, effectively avoiding the deflection and rerouting issues typically encountered in bufferless NoCs. It was shown that segmented buses such as NeuSB [balaji2022NeuSB] outperforms NoCs in both latency and energy consumption for running SNN applications. Nevertheless, NeuSB lacks runtime switch controls and thus does not truly support running different applications on the same hardware. In this work, we focus on segmented ladder bus [huynh2024adiona], the latest dynamic segmented bus interconnect that can offer runtime configuration for routing and scheduling network traffic.
@@ 2 Background and Related Works
This section presents the requisite background to comprehend the segmented ladder bus [huynh2024adiona], the primary interconnect architecture targeted by our mapping and scheduling methodologies. Furthermore, we review related works on the mapping of SNNs to neuromorphic hardware architectures, emphasizing the associated challenges and recent innovations in this domain.
@@ 2.1 Segmented Ladder Bus
A segmented ladder bus is governed globally by a compile time software framework and locally use coarse-grain runtime hardware controllers in order to enable dynamic configuration of switches. These local controllers encompass predefined switch change scenarios and facilitate the transmission of control signals to the switches accordingly. By issuing the control signals from the controllers with the correct timing, the switches can be used to create paths that support multiple simultaneous connections inside the network, matching application requirements. The controllers are implemented in a distributed manner, where each controller is responsible for managing the switches within its vicinity, similar to a distributed loop controller [raghavan2008distributed]. As a result, in a large-scale network, numerous controllers will be deployed to effectively govern the switches throughout the system.
@@
Compare to NoCs, the segmented ladder bus sacrifices general-purpose routing flexibility in exchange for compile-time optimized, energy-efficient communication tailored to the sparse and bursty traffic spiking trains in neuromorphic workloads. While NoCs offer higher runtime flexibility and better fault tolerance due to their adaptive routing and buffering, they also incur significant static and dynamic energy overheads and complexity, which are often unnecessary for SNNs. The ladder bus, being bufferless and pre-scheduled, can operate at lower power and area cost while meeting the timing and bandwidth requirements of neuromorphic applications. Fault tolerance in our approach can be addressed at the compile-time mapping level by reassigning communication paths around faulty switches or tiles.
Although the segmented ladder bus is bufferless and lacks intermediate synchronization points, this does not hinder its suitability for neuromorphic systems. In this domain, precise timing closure is not a strict requirement, as these systems are event-driven and tolerant to timing variations [serrano2013stdp]. As long as spike event ordering and causality are preserved, variations in communication latency across longer paths do not compromise application correctness. This is because spike trains typically have inter-spike timing distances in the order of milliseconds or higher. Therefore, while signal delays may vary slightly across long paths, potentially reaching µsecond levels, the impact on application level correctness is negligible due to the above-mentioned millisecond-scale timing requirements of neuromorphic workloads. This temporal flexibility makes the segmented ladder bus particularly well-suited for energy-efficient, pre-scheduled interconnects that do not require cycle-level synchronization across the chip.
A more comprehensive understanding of the advantages offered by segmented ladder bus can be gained through an examination of Figure 2. The cluster communication graph is presented on the left with each cluster denoted by a unique identifier from C0 to C8. In order to be deployed for execution, each cluster is sized so that it fits on a single hardware tile (neuromorphic core). With a suitable mapping (e.g., C0 to T4, C1 to T1, etc.), multiple communication paths can be facilitated simultaneously: Tile 0 sends spikes to Tiles 5 and 6, Tile 1 sends to Tile 2, and Tile 7 to Tiles 3 and 8.
@@ 2.2 Related Works
One prominent method is SpiNeMap [spinemap], which focuses on optimizing the mapping of spiking neural networks (SNNs) onto neuromorphic hardware. SpiNeMap uses multi-objective optimization to balance energy, latency, and inter-spike interval based on SNN-specific communication patterns. Similarly, SNEAP [li2020sneap] employs heuristics such as Simulated Annealing, Particle Swarm Optimization, and Tabu Search, but its dependence on static NoC topologies limits scalability and adaptability under dynamic workloads.
Jin et al. [jin2023mapping] propose a heuristic approach combining Hilbert curve-based placement with a Force Directed refinement, leveraging SNN sparsity to reduce communication overhead. Despite its effectiveness, the heuristic nature of this algorithm may not always guarantee optimal mappings, especially in larger and more complex networks.
NeuMap [neumap] introduces a machine learning-based mapper that adapts to workload variations using predictions from trained models. Though promising in adaptability, it incurs training overhead and depends on extensive historical data, which may hinder real-time deployment.
@@ 3 Proposed Solutions
In our previous work [mustafazade2024clustering], we have described the SNN application partitioning process in more details. This allocates neurons to memristive crossbar clusters while maximizing the crossbar utilization and minimizing inter-cluster communication. Although such process is essential, it works independently of the types of interconnect hardware. In particular, the algorithms take only the spike traffic information from the application as input and do not account for the architectural constraints or design opportunities introduced by the segmented ladder bus or other specific interconnects. For this reason, it is not the focus of this paper.
Once an SNN application has been partitioned into clusters, we need to: map the clusters to actual hardware tiles; schedule spike traffic; and route the scheduled traffic. Each of these steps can be optimized to improve the hardware utilization of the network, guarantee performance of the application, and reduce energy consumption. Figure 3 shows our proposed solution processes integrated inside our evaluation framework. While numerous state-of-the-art network simulators provide support for Network on Chip architectures, none currently offer built-in capabilities for simulating segmented bus architectures. Thus, we developed a custom network simulator tailored to segmented busesbased on Noxim [catania2015noxim]. Noxim was chosen as the foundation for our simulator due to its open-source, extensible codebase, SystemC-based [systemC] cycle-accurate simulation engine, and lightweight, scalable design.
@@ 3.1 Cluster Mapping
After an SNN application has been partitioned into clusters, it is important to map these clusters into actual hardware tiles. The connectivity between the clusters can be represented as a directed graph $G_{\mathit{CSNN}}=(C,L)$} consisting of $C$ clusters and $L$ links between the clusters. A link $L_{i,j}\in L$} is a channel connecting cluster $C_{i}$ with cluster $C_{j}$, where $C_{i}$, $C_{j}\in C$. Each link $L_{i,j}$ has a weight $s_{i,j}$ that represents the number of spikes communicated on $L_{i,j}$. A spike link activity indicator $z_{i,j}=1$ if $s_{i,j}>0$ and $s_{i,j}=0$ otherwise. The cluster mapping problem can be modeled as an injective function $\phi$ that place the neuron clusters $C$ into hardware tiles $T$.
Different mappings will have different impacts on energy consumption, latency, and the number of crossed communication paths. In our segmented ladder bus design, communication energy consumption is proportional to the length of the communication paths and the number of switch configurations that need to be changed. Furthermore, unlike NoCs, the bufferless nature of the switches in segmented bus causes simultaneous intersecting communication paths to result in either packet drops or additional delays at the source nodes to avoid packet loss. Thus, we propose two different mapping cost functions specifically adapted to ladder segmented bus to minimize either dynamic energy consumption or number of crossed paths. We focus on dynamic energy because, for any given application, the same hardware is used across all mapping, scheduling, and routing configurations. As a result, static (leakage) power remains largely constant and does not impact the relative comparison between configurations. Including static power would only introduce a uniform offset to the total energy values. Therefore, dynamic energy more accurately reflects the impact of interconnect activity and the effectiveness of different mapping strategies. They are computed as follows.
@@
Given that cluster mapping is a well-established NP-hard problem [gary1979computers], a practical solution is implemented using a Hill Climbing approach [selman2006hill]. This metaheuristic provides an efficient and practical way to explored the vast solution space while balancing computational cost and solution quality. We employ a specific variant, called steepest ascent Hill Climbing, to enhance the search process. The algorithm starts with a random mapping solution and compares the cost of that solution with all neighboring solutions. A neighbor is defined as a mapping obtained by swapping the positions of two nodes in the current mapping. This ensures that the search space of the algorithm is limited to only valid mappings, which improve the efficiency of the search. The algorithm moves the solution to the neighbor with the most improved cost. The search continues until there is no neighbor with improved cost. The local search ends here, with the algorithm keeping record of this solution. It is important to note that Hill Climbing may converge to a local minima. To mitigate this, when the search terminates due to the lack of better-cost neighbors, we incorporate a number of random swaps and re-initiate the search procedure. Ultimately, the algorithm concludes its execution after running a predetermined number of random perturbations.
Input: $G_{\mathit{CSNN}}$, $\eta$, II
@@ 6  $k=0$;
7  while improvedimproved and $k<I$ do /* while cost is improved and under the number of iteration limit */
@@ 3.2 Spike Traffic Scheduling
With the cluster mapping acquired from algorithm 1 and the spike traffic from CARLSim [carlsim6] or SNNTorch [snntorch], we can obtain the exact timing of the spike traffic among the hardware tiles. For any given time with multiple simultaneous communication links, the scheduler would need to find groups of links for which no two links in any group intersect, while minimizing the number of groups to prevent unnecessary delays.
This problem is essentially a bin packing problem with conflicts (BPPC) [ekici2021bin], which is NP-complete. Therefore, instead of seeking an optimal schedule, we develop a custom heuristic to get a solution within a practical computational limit. An example of a spike traffic schedule can be found in Figure 5.
@@ 8        $crossed=False$;
10           if $check\_cross(G_{ladder},L,L_{G})$ then /* check whether the current link L crosses with $L_{G}$ */
@@ 29 $delay=0$;
    For each link $l$, the algorithm checks existing groups to see if the link can be added without conflict. For each group, the algorithm checks if the new link crosses with any existing link in the group. This involves iterating over each link in the group, and for each pair of links, the check_cross function is called, which can be simplified to take constant time $O(1)$ due to the defined and regular topology of the network. In the worst case, each link will be compared to all other L1L-1 links, giving $O(L^{2})$ comparisons.
@@ 3.3 Path Routing
Our solution to the path routing problem primarily adopts a greedy algorithm, supplemented by A* [duchovn2014path] sub-processes for efficient shortest path finding. We sort all the simultaneous communication links $CLinks$ according to the following order: we prioritize all the links with source and destination tiles on the same bus lane first, followed by sorting based on the orientation of the links, and finally by the number of spikes on each link. The reasoning of this sort is three folds: Firstly, we reduce the number of potential crossing between links by routing the links with source and destination tiles on the same bus lane and the shorter distance ones first. This is because the shortest path for tiles on the same bus lane is unique, while there can be many paths with the same length for tiles on different bus lanes. Secondly, links with the same orientation can be easily routed one by one so that they do not cross each other. Thirdly, by prioritizing to route the links with more spikes, the paths for these links will be shorter, which means lowering the number of spikes that need to travel long distances, which in turn helps in reducing the energy consumption in the network. We then represent the segmented ladder bus interconnect as an undirected graph $G_{ladder}$ with tiles and switches as vertices and bus wires as edges. The initial weights of all vertices and edges in the graph are set to 1. The algorithm processes the sorted communication links sequentially, and uses a modified A* algorithm to determine the shortest weighted path for each specific link. After determining the shortest path for the selected link, the weights of all vertices are increased proportionally to the number of spikes communicated on this particular link. Subsequently, the routed link is removed from the list of communication links. The algorithm continues the process iteratively, selecting and routing the most communicated links one by one, until all paths are effectively routed.
Input: $G_{ladder}$, CLinksCLinks
Output: RoutedListRoutedList